OpenStudy (anonymous):

Help would be appreciated! Which of the following are vertical asymptotes of the function y=2ct(3x)+4? (Check all that apply) A) x=0 B) x=2pi C) x= +/- pi/2 D)x= pi/3

OpenStudy (johnweldon1993):

Is that \[\large y = 2cot(3x) + 4\] ?

OpenStudy (anonymous):

Yes! @johnweldon1993

OpenStudy (johnweldon1993):

Hmm well since vertical asymptotes occur when the function approaches infinity, and that happens when a denominator = 0 We know that \(\large cot(3x) = \frac{1}{tan(3x)}\) We ask ourselves when tan(3x) = 0? Well to my knowledge, that only occurs at tan(0) so it looks like x = 0 is your vertical asymptote here

OpenStudy (anonymous):

Thank you! Would that be the only one or is there another possibility? @johnweldon1993

OpenStudy (johnweldon1993):

Well from what the graph looks like, it looks as if there is a repeating vertical asymptote every \(\large \frac{\pi}{3}\) But lets check that

OpenStudy (johnweldon1993):

\[\large y = \frac{2}{tan(3x)} + 4\] \[\large y = \frac{2cos(3x)}{sin(3x)} + 4\] So that is where we see that at x = 0 we approach infinity...but now lets check \(\large \frac{\pi}{3}\) \[\large y = \frac{2cos(\frac{3\pi}{3})}{(sin\frac{3\pi}{3})} + 4\] \[\large y = \frac{2cos(\pi)}{sin(\pi)} + 4\] Ahh and also \(\large sin(\pi) = 0\) so yes also one every \(\large \frac{\pi}{3}\)

OpenStudy (johnweldon1993):

haha now that I look, another one of your answers is correct!

OpenStudy (johnweldon1993):

So as you can see...the main thing I look for is when the denominator = 0 The denominator here is \(\large sin(3x)\) So whenever that = 0 we have a vertical asymptote

OpenStudy (johnweldon1993):

We know \(\large sin(0) = 0\) \[\large sin(\pi) = 0\] \[\large sin(2\pi) = 0\] and also any multiple of \(\large \pi\) = 0 when it comes to sin...so It looks like A, B, and D are ALL correct here

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (johnweldon1993):

No problem!